Forces and Work (c6p4) Two forces parallel to the x axis do 16.2 J of work on a small tray while moving it 16.2 m in the x direction across a gym floor. One of the forces has a value of +3.58 N in thex direction. What is the other force? 3.58 N Submit Answer Incorrect. Tries 3/10 Previous Tries
Answer #1

Work done, W = (F1 + F2)*d

F1 + F2 = W/d

F2 = W/d - F1

Therefore, F2 = 16.2/16.2 - 3.58

F2 = -3.58 N